Estimate Fresnel distance for slit of size 1 cm for wavelength 500nm

Grade:12

1 Answers

Arun
6 years ago
Fresnel’s distance (ZF) is the distance for which the ray optics is a good approximation. It is given by the relation,
Zf = a²/lambda = 1*10^-2 *1*10^-2 /500*10^-9
= 10^-4/5*10^-7
= 1000/5 = 200 m
 
Regards
Arun (askIITians forum expert)